# Deriving a GENERIC system from a Hamiltonian system

### v2

@@ Abstract
Already in the finite-energy case (zero-temperature heat bath) we obtain the so-called GENERIC structure (General Equations for Non-Equilibrium Reversible Irreversibe Coupling), with conserved energy, nondecreasing entropy, a new Poisson structure, and an Onsager operator describing the dissipation. However, their origin is not obvious at this stage. After extending the system in a natural way to the case of positive temperature, giving a heat bath with infinite energy, the compression property leads to an exact multivariate Ornstein-Uhlenbeck process that drives the rest of the system. Thus, we are able to identify a conserved energy, an entropy, and an Onsager operator (involving the Green-Kubo formalism) which indeed provide a GENERIC structure for the macroscopic system.
@@ 1.1 Coarse-graining, GENERIC, and compressions
Coarse-graining is the process by which less detailed (‘coarse-grained’) models of physical systems are derived from more detailed ones (‘fine-grained’; see, e.g., [E11, Kue15]). One methodology for coarse-graining is the Mori-Zwanzig formalism, which allows to give macroscopic descriptions of microscopic systems such as those arising in molecular dynamics. In particular, this allows one to study the emergence of macroscopic dissipation from microscopically reversible Hamiltonian systems. In this paper we study a classical example of this type [JaP98a, Zwa01a].
GENERIC [Mor86, GrÖ97, ÖtG97] describes a class of thermodynamic evolution equations. If an evolution equation can be put in GENERIC form, then it is automatically thermodynamically consistent. In particular, it satisfies conservation of an energy and monotonicity of an entropy. In GENERIC, this thermodynamic consistency is implemented by combining a symplectic operator acting on the energy and a semidefinite operator acting on the entropy. In addition, the book [Ö05] describes a Mori-Zwanzig-type derivation of equations of GENERIC type from microscopic Hamiltonian systems with a clear separation of time scales. In this paper we derive a GENERIC form for the coarse-grained evolution, and this allows us to give an interpretation for the various components of the GENERIC structure.
@@ 1.2 Sketch of the microscopic Hamiltonian system
The starting point of this paper, the ‘microscopic system’, resembles well-known models in the literature (see, e.g., [JaP98a, KS02a, Rey06]). It couples two separate Hamiltonian systems: System A is finite-dimensional, and System B is infinite-dimensional and is interpreted as a heat bath. System A may be nonlinear, while System B and the coupling are linear. This is illustrated in Figure 1.1. The whole system is Hamiltonian, and in the interesting cases, the system also is reversible in time. The evolution is deterministic, while the initial data for the heat bath may be random, as we discuss below.
@@
${\mathcal{H}}_{\mathcal{C}}=\left({\mathbb{A}}z\middle|{\mathbb{P}}\eta\right)_{\mathbf{H}},$ (1.1)
@@
${\mathcal{H}}_{\mathcal{C}}=\left({\mathsf{C}}z\middle|\eta\right)_{\mathbf{H}}.$ (1.2)
@@
and the full Hamiltonian system is denoted by the triple $({\mathbf{Z}}{\times}{\mathbf{H}},{\mathcal{H}}_{\mathrm{total}},{\mathbb{J}}_{\mathcal{A}}{\oplus}{\mathbb{J}}_{\mathcal{B}})$.
@@
In particular, the unitary group $(\mathrm{e}^{t{\mathbb{J}}_{\mathcal{B}}})_{t\in{\mathbb{R}}}$ is the group of shifts ${\mathbb{S}}_{t}f:=f(\,\cdot\,{-}t)$. For the coupling we assume that only the state variable $q$ (but not the momentum $p$) couples with the heat bath, i.e. ${\mathcal{H}}_{\mathcal{C}}(q,p,\eta)=\left(\mathsf{C}q\middle|\eta\right)$ for a linear bounded operator $\mathsf{C}\colon{\mathbb{R}}^{n}\to{\mathbf{H}}$.
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\begin{array}[]{c}q\\ p\\ \eta\end{array}\right)=\left(\begin{array}[]{ccc}0&{\mathbb{I}}_{n\times n}\!\!&0\\ \!\!{-}{\mathbb{I}}_{n\times n}\!\!&0&0\\ 0&0&\!\!{-}\partial_{x}\!\end{array}\right)\left(\begin{array}[]{c}{\mathrm{D}}V(q)+{\mathrm{D}}_{q}{\mathcal{H}}_{\mathcal{C}}(q,\eta)\\ p\\ \eta+{\mathrm{D}}_{\eta}{\mathcal{H}}_{\mathcal{C}}(q,\eta)\end{array}\right)=\left(\begin{array}[]{c}p\\ \!\!-{\mathrm{D}}V(q)-\mathsf{C}^{*}\eta\!\!\\[3.00003pt] -\partial_{x}(\eta+\mathsf{C}q)\end{array}\right).$
The first two equations are ODEs in ${\mathbb{R}}^{d}$, whereas the third equation is a transport equation in ${\mathrm{L}}^{2}({\mathbb{R}};{\mathbb{R}}^{3})$ that has to be interpreted in a suitable weak or mild form, if we want to treat general initial conditions (see the abstract construction in Theorem 2.5).
@@ 1.3 The compression property
Having the explicit representation of ${\mathbf{W}}$, the coupling $\mathsf{C}\colon{\mathbb{R}}^{n}\to{\mathbf{W}}\subset{\mathbf{H}}$ can be specified as $\mathsf{C}q=(\mathsf{c}^{1}{\cdot}q)f_{1}+(\mathsf{c}^{2}{\cdot}q)f_{2}+(\mathsf{c}^{3}{\cdot}q)f_{3}$, where $\mathsf{c}^{1},\mathsf{c}^{2},\mathsf{c}^{3}\in{\mathbb{R}}^{n}$ are given coupling vectors. Hence, the coupling energy reads
${\mathcal{H}}_{\mathcal{C}}(q,p,\eta)=\left(\mathsf{C}q\middle|{\mathbb{P}}\eta\right)_{\mathbf{H}}=\sum_{j=1}^{3}(\mathsf{c}^{j}{\cdot}q)\int_{-\infty}^{0}\!\!f_{j}(x){\cdot}\eta(x)\;\!\mathrm{d}x,$
@@ 1.4 Coarse-graining, part 1: Deterministic initial data
To describe this result precisely we first define the coarse-graining map,
@@
In the notation introduced for GENERIC, we have ${\mathbf{Y}}:={\mathbf{Z}}{\times}{\mathbf{W}}{\times}{\mathbb{R}}$. Note how this map $\pi_{z,w,e}$ preserves the System-A state variable $z$ but projects the heat-bath variable $\eta$ onto the finite-dimensional subspace ${\mathbf{W}}$. The third variable $e$ allows us to recover an important piece of information from the missing part $\eta-w=\eta-{\mathbb{P}}\eta$, namely its total energy $\tfrac{1}{2}\|\eta-w\|_{\mathbf{H}}^{2}=\tfrac{1}{2}\|\eta\|_{\mathbf{H}}^{2}-\tfrac{1}{2}\|w\|_{\mathbf{H}}^{2}$. The variable $e$ is only used to keep track of the energy exchanged with the heat bath. Already for $(z,w)=(z,{\mathbb{P}})$ we obtain the closed system of ODEs
@@ 1.6 Coarse-graining, part 3: Random initial data
=${}_{\mathrm{GEN}}{}_{\mathrm{GEN}}^{*}=2{\mathbb{K}}_{\mathrm{GEN}}.$
@@ 1.7 Outline of this paper
In Section 4 we describe the GENERIC structure in detail and establish some properties of so-called GENERIC SDEs. By taking into account the energy flowing in and out of the heat bath, as described in in Sections 1.4 and 1.6, we can formulate the equations obtained in Section 3.2 as a GENERIC system, either deterministic or stochastic, and identify the components. This also gives insight in how the GENERIC restrictions arise in the coarse-graining process.
@@ Acknowledgements
The authors wish to thank Noé Corneille, Greg Pavliotis, Artur Stephan, and the ‘Wednesday morning session’ at Eindhoven University of Techology for valuable input during the preparation of this manuscript.
@@ 2 The microscopic system
The microscopic setup of this paper is a well-known model (see, e.g., [JaP98a, KS02a, Rey06]). As mentioned above, two separate Hamiltonian systems are coupled, where one is finite-dimensional (‘System A’), and the other is infinite-dimensional and is interpreted as a heat bath (‘System B’). System A may be nonlinear, while System B and the coupling will be linear. This is illustrated in Figure 1.1.
@@ 2.1 Finite energy, deterministic evolution
In the context of Hilbert spaces, there is a choice to consider either the Fréchet derivative, which is an element of the dual space, or the ‘gradient’, which is the Riesz representative of the Fréchet derivative and therefore an element of the primal space. In this paper we always choose the second interpretation, as the Riesz representatitive of the Fréchet derivative. With this choice, we have for instance for all $\eta\in{\mathbf{H}}$ and $z\in{\mathbf{Z}}$,
@@ 2.2 Positive temperature: The heat bath
The evolution operator $\mathrm{e}^{t{\mathbb{J}}_{\mathcal{B}}}$, that describes the heat-bath unitary evolution in the smaller space ${\mathbf{H}}$, can similarly be extended $\gamma_{\beta}$-a.e. to a operators $\widehat{S}_{t}$ on ${\mathbf{X}}$. Without loss of generality we make the following simplifying assumption (see Section C.2 for a discussion):
@@
    Choose the probability space $({\mathbf{X}},{\mathcal{B}}({\mathbf{X}})_{\gamma_{\beta}},\gamma_{\beta})$, where ${\mathcal{B}}({\mathbf{X}})_{\gamma_{\beta}}$ is the $\gamma_{\beta}$-completion of the Borel $\sigma$-algebra ${\mathcal{B}}({\mathbf{X}})$. If $\eta_{0}$ is drawn from ${\mathbf{X}}$ with law $\gamma_{\beta}$, then both $Y=Y(\eta_{0})$ and $Y=Y(\widehat{\mathbb{Q}}\eta_{0})$ are measurable ${\mathbf{W}}$-valued stochastic processes with sample paths in $L^{2}_{\mathrm{loc}}({\mathbb{R}};{\mathbf{W}})$; in addition $Y(\eta_{0})$ is stationary.
@@
The additional characterization of $Y(\widehat{\mathbb{Q}}\eta_{0})$ will be useful below, when we solve the initial-value problem with a given value of $w_{0}:=\widehat{\mathbb{P}}\eta_{0}$.
@@ 2.3 Well-posedness of the stochastic process on the product space
With the characterization of the process $Y$ by Lemma 2.4 we can now construct the solution of the coupled microscopic system on the product space ${\mathbf{Z}}{\times}{\mathbf{H}}$ (deterministic and with finite energy) or ${\mathbf{Z}}{\times}{\mathbf{X}}$ (random and with infinite energy).
@@
where $Y=Y(\eta_{0})$. For $(z_{0},\eta_{0})\in{\mathbf{Z}}{\times}{\mathbf{H}}$, the process $t\mapsto Y_{t}(\eta_{0})$ is continuous with values in ${\mathbf{W}}$ by Lemma 2.4, and a standard ODE proof based on contraction for small times gives short-time existence and uniqueness of a solution of (2.10) with the specified regularity. The global Lipschitz bound on ${\mathrm{D}}{\mathcal{H}}_{\mathcal{A}}$ implies that the time of existence does not depend on the initial datum, and we obtain existence globally in time. Equation (2.8a) can then be recovered from (2.10) by differentiating in time, and equation (2.8b) can be used to reconstruct $\eta$. This proves the part 1 of the theorem.
@@
For our model discussed in Examples 1.1 and 1.3 we can explicitly construct the larger Hilbert space ${\mathbf{X}}\supset{\mathbf{H}}={\mathrm{L}}^{2}({\mathbb{R}};{\mathbb{R}}^{3})={\mathrm{L}}^{2}({\mathbb{R}})^{3}$. In fact, this can be done independently for each of the three components. We follow the classical construction in [Rey06] and choose
@@
Here $X$ contains distributions $T$ that can be decomposed into $T=T_{1}+T_{2}$ with $T_{2}\in{\mathrm{H}}^{-2}(R)$ and $T_{2}\in{\mathrm{L}}^{2}_{-2}({\mathbb{R}})$, i.e. $x\mapsto(1{+}x^{2})^{-1}T_{2}(x)\in{\mathrm{L}}^{2}({\mathbb{R}})$. Clearly, the shift operator $\widehat{S}_{t}:T\mapsto T(\!\;\cdot\;\!{-}t)$ generates a continuous semigroup that is the extension of the unitary shift group $(\mathrm{e}^{t\partial_{x}})_{t\in{\mathbb{R}}}$ on ${\mathrm{L}}^{2}({\mathbb{R}})$.

### v3

@@ Abstract
Already in the finite-energy case (zero-temperature heat bath) we obtain the so-called GENERIC structure (General Equation for Non-Equilibrium Reversible Irreversible Coupling), with conserved energy, nondecreasing entropy, a new Poisson structure, and an Onsager operator describing the dissipation. However, their origin is not obvious at this stage. After extending the system in a natural way to the case of positive temperature, giving a heat bath with infinite energy, the compression property leads to an exact multivariate Ornstein-Uhlenbeck process that drives the rest of the system. Thus, we are able to identify a conserved energy, an entropy, and an Onsager operator (involving the Green-Kubo formalism) which indeed provide a GENERIC structure for the macroscopic system.
@@ 1.1 Coarse-graining, GENERIC, and compressions
Coarse-graining is the process by which less detailed (‘coarse-grained’) models of physical systems are derived from more detailed ones (‘fine-grained’; see, e.g., [E11, Kue15]). One methodology for coarse-graining is the Mori-Zwanzig formalism, which allows to give macroscopic descriptions of microscopic systems such as those arising in molecular dynamics. In particular, this allows one to study the emergence of macroscopic dissipation from microscopically reversible Hamiltonian systems. In this paper we study a classical example of this type; see [JaP98, Zwa73] and also the monograph [Zwa01] for related treatments.
GENERIC describes a class of thermodynamic evolution equations that is formulated in terms on an energy and an entropy functional and two geometric structures for the Hamiltonian and the dissipative dynamics. The origins of such systems goes back to [Grm84, Grm85] and the metriplectic theory developed in [Mor84, Mor86, Mor09], whereas the name General Equation for Non-Equilibrium Reversible Irreversible Coupling and the acronym GENERIC were introduced in [GrÖ97, ÖtG97]. The GENERIC framework provides a general way to write down thermodynamically consistent equations for the coupling of two ideal systems, namely a reversible (or Hamiltonian) system and an irreversible (or dissipative) system. GENERIC systems form a subclass of all thermodynamically consistent systems that allow for a better understanding of the interaction between the two different phenomena of reversible and irreversible dynamics. In particular, it will help us to see how dissipative systems can be obtained via coarse graining of a purely Hamiltonian systems.
If an evolution equation can be put in GENERIC form, then it is automatically thermodynamically consistent. In particular, it satisfies conservation of an energy and monotonicity of an entropy. In GENERIC, this thermodynamic consistency is implemented by combining a symplectic operator acting on the energy and a semidefinite operator acting on the entropy. In addition, the book [Ött05] describes a Mori-Zwanzig-type derivation of equations of GENERIC type from microscopic Hamiltonian systems with a clear separation of time scales. In this paper we derive a GENERIC form for the coarse-grained evolution, and this allows us to give an interpretation for the various components of the GENERIC structure.
@@ 1.2 Sketch of the microscopic Hamiltonian system
The starting point of this paper, the ‘microscopic system’, resembles well-known models in the literature (see, e.g., [JaP98, KS02, Rey06]). It couples two separate Hamiltonian systems: System A is finite-dimensional, and System B is infinite-dimensional and is interpreted as a heat bath. System A may be nonlinear, while System B and the coupling are linear. This is illustrated in Figure 1.1. The whole system is Hamiltonian, and in the interesting cases, the system also is reversible in time. The evolution is deterministic, while the initial data for the heat bath may be random, as we discuss below.
@@
${\mathcal{H}}_{\mathcal{C}}(z,\eta)=\left({\mathbb{A}}z\middle|{\mathbb{P}}\eta\right)_{\mathbf{H}},$ (1.1)
@@
${\mathcal{H}}_{\mathcal{C}}(z,\eta)=\left({\mathsf{C}}z\middle|\eta\right)_{\mathbf{H}}.$ (1.2)
@@
The bilinearity of (1.2) is assumed for simplicity; we expect that similar results can be proved also for functionals that are nonlinear in $z$ but linear in $\eta$, e.g., the model studied in [Zwa73].
@@
and the full Hamiltonian system is denoted by the triple $({\mathbf{Z}}{\times}{\mathbf{H}},{\mathcal{H}}_{\mathrm{total}},{\mathbb{J}}_{\mathcal{A}}{\oplus}{\mathbb{J}}_{\mathcal{B}})$. The aim of this paper is to coarse-grain the coupled system (1.4) to a system where the macroscopic observable $z\in{\mathbf{Z}}$ is maintained, but the infinite-dimensional degree of freedom $\eta\in{\mathbf{H}}$ is reduced to a finite-dimensional coupling variable $w$. The so-called compression property (see Assumption 1.3 below) will imply that despite the loss of information the resulting system in $(z,w)\in{\mathbf{Z}}{\times}{\mathbf{W}}$ still is Markovian.
@@
In particular, the unitary group $(\mathrm{e}^{t{\mathbb{J}}_{\mathcal{B}}})_{t\in{\mathbb{R}}}$ is the group of shifts ${\mathbb{S}}_{t}f:=f(\,\cdot\,{-}t)$. For simplicity we assume in this example that coupling to the heat bath occurs only via the state variable $q$ (but not the momentum $p$), i.e. ${\mathcal{H}}_{\mathcal{C}}(q,p,\eta)=\left({\mathsf{K}}q\middle|\eta\right)$ for a linear bounded operator ${\mathsf{K}}\colon{\mathbb{R}}^{n}\to{\mathbf{H}}$.
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\begin{array}[]{c}q\\ p\\ \eta\end{array}\right)=\left(\begin{array}[]{ccc}0&{\mathbb{I}}_{n\times n}\!\!&0\\ \!\!{-}{\mathbb{I}}_{n\times n}\!\!&0&0\\ 0&0&\!\!{-}\partial_{x}\!\end{array}\right)\left(\begin{array}[]{c}{\mathrm{D}}V(q)+{\mathrm{D}}_{q}{\mathcal{H}}_{\mathcal{C}}(q,\eta)\\ p\\ \eta+{\mathrm{D}}_{\eta}{\mathcal{H}}_{\mathcal{C}}(q,\eta)\end{array}\right)=\left(\begin{array}[]{c}p\\ \!\!-{\mathrm{D}}V(q)-{\mathsf{K}}^{*}\eta\!\!\\[3.00003pt] -\partial_{x}(\eta+{\mathsf{K}}q)\end{array}\right).$
The first two equations are ODEs in ${\mathbb{R}}^{n}$, whereas the third equation is a transport equation in ${\mathrm{L}}^{2}({\mathbb{R}};{\mathbb{R}}^{3})$ that has to be interpreted in a suitable weak or mild form, if we want to treat general initial conditions (see the abstract construction in Theorem 2.5).
@@ 1.3 The compression property
Remark 1.3 (The compression property replaces ergodicity).
In finite-dimensional Hamiltonian systems, dissipative behavior is often associated with ergodic properties. In the context of the infinite-dimensional Hamiltonian heat bath of this paper, the Compression Property 1.3 mimics the property of ergodic systems. In the rest of this paper, we indeed show that the resulting system has thermodynamically consistent dissipative behavior.
@@
Having the explicit representation of ${\mathbf{W}}$, the coupling ${\mathsf{K}}\colon{\mathbb{R}}^{n}\to{\mathbf{W}}\subset{\mathbf{H}}$ can be specified as ${\mathsf{K}}q=(\mathsf{c}^{1}{\cdot}q)f_{1}+(\mathsf{c}^{2}{\cdot}q)f_{2}+(\mathsf{c}^{3}{\cdot}q)f_{3}$, where $\mathsf{c}^{1},\mathsf{c}^{2},\mathsf{c}^{3}\in{\mathbb{R}}^{n}$ are given coupling vectors. Hence, the coupling energy reads
${\mathcal{H}}_{\mathcal{C}}(q,p,\eta)=\left({\mathsf{K}}q\middle|{\mathbb{P}}\eta\right)_{\mathbf{H}}=\sum_{j=1}^{3}(\mathsf{c}^{j}{\cdot}q)\int_{-\infty}^{0}\!\!f_{j}(x){\cdot}\eta(x)\;\!\mathrm{d}x,$
@@ 1.4 Coarse-graining, part 1: Deterministic initial data
To describe this result precisely we first define the coarse-graining map $\pi_{z,w,e}$,
@@
Thus, we apply the coarse-graining map to the coupled system $\big{(}{\mathbf{Z}}{\times}{\mathbf{H}},{\mathcal{H}}_{\mathrm{total}},{\mathbb{J}}_{\mathcal{A}}{\oplus}{\mathbb{J}}_{\mathcal{B}}\big{)}$, which is an infinite-dimensional Hamiltonian system, and obtain a coarse-grained system on the reduced state space ${\mathbf{Y}}:={\mathbf{Z}}{\times}{\mathbf{W}}{\times}{\mathbb{R}}$. This is done by projecting out infinitely many degrees of freedom in ${\mathbf{H}}$ and keeping only the new observables $w={\mathbb{P}}\eta$ in the finite-dimensional subspace ${\mathbf{W}}$ of ${\mathbf{H}}$. Note that System A is considered already as a ‘coarse-grained’ system, which is preserved by $\pi_{z,w,e}$, and coupled to the new observable $w\in{\mathbf{W}}$. The third variable $e$ allows us to recover an important piece of information from the missing part $\eta-w=\eta-{\mathbb{P}}\eta$, namely its total energy $\tfrac{1}{2}\|\eta-w\|_{\mathbf{H}}^{2}=\tfrac{1}{2}\|\eta\|_{\mathbf{H}}^{2}-\tfrac{1}{2}\|w\|_{\mathbf{H}}^{2}$. The variable $e$ is only used to keep track of the energy exchanged with the heat bath. Already for $(z,w)=(z,{\mathbb{P}}\eta)$ we obtain the closed system of ODEs
@@ 1.6 Coarse-graining, part 3: Random initial data
=${}_{\mathrm{GEN}}{}_{\mathrm{GEN}}^{*}=\frac{2}{\beta}{\mathbb{K}}_{\mathrm{GEN}}.$
@@ 1.7 Outline of this paper
In Section 4 we describe the GENERIC structure in detail and establish some properties of so-called GENERIC SDEs. By taking into account the energy flowing in and out of the heat bath, as described in Sections 1.4 and 1.6, we can formulate the equations obtained in Section 3.2 as a GENERIC system, either deterministic or stochastic, and identify the components. This also gives insight in how the GENERIC restrictions arise in the coarse-graining process.
@@ Acknowledgements
The authors wish to thank Noé Corneille, Greg Pavliotis, Artur Stephan, and the ‘Wednesday morning session’ at Eindhoven University of Techology for valuable input during the preparation of this manuscript. The research of A.M. has been partially funded by Deutsche Forschungsgemeinschaft (DFG) through grant CRC 1114 “Scaling Cascades in Complex Systems” (project no. 235221301), Subproject C05 Effective Models for Materials and Interfaces with Multiple Scales.
@@ 2 The microscopic system
The microscopic setup of this paper is a well-known model (see, e.g., [JaP98, KS02, Rey06]). As mentioned above, two separate Hamiltonian systems are coupled, where one is finite-dimensional (‘System A’), and the other is infinite-dimensional and is interpreted as a heat bath (‘System B’). System A may be nonlinear, while System B and the coupling will be linear. This is illustrated in Figure 1.1.
@@ 2.1 Finite energy, deterministic evolution
In the context of Hilbert spaces, there is a choice to consider either the Fréchet derivative, which is an element of the dual space, or the ‘gradient’, which is the Riesz representative of the Fréchet derivative and therefore an element of the primal space. In this paper we always choose the second interpretation, as the Riesz representative of the Fréchet derivative. With this choice, we have for instance for all $\eta\in{\mathbf{H}}$ and $z\in{\mathbf{Z}}$,
@@ 2.2 Positive temperature: The heat bath
The evolution operator $\mathrm{e}^{t{\mathbb{J}}_{\mathcal{B}}}$, that describes the heat-bath unitary evolution in the smaller space ${\mathbf{H}}$, can similarly be extended $\gamma_{\beta}$-a.e. to an operator $\widehat{S}_{t}$ on ${\mathbf{X}}$. Without loss of generality we make the following simplifying assumption (see Section C.2 for a discussion):
@@
    Choose the probability space $({\mathbf{X}},{\mathcal{B}}({\mathbf{X}})_{\gamma_{\beta}},\gamma_{\beta})$, where ${\mathcal{B}}({\mathbf{X}})_{\gamma_{\beta}}$ is the $\gamma_{\beta}$-completion of the Borel $\sigma$-algebra ${\mathcal{B}}({\mathbf{X}})$. If $\eta_{0}$ is drawn from ${\mathbf{X}}$ with law $\gamma_{\beta}$, then both ${Y_{t}}={Y_{t}}(\eta_{0})$ and ${Y_{t}}={Y_{t}}(\widehat{\mathbb{Q}}\eta_{0})$ are measurable ${\mathbf{W}}$-valued stochastic processes with sample paths in $L^{2}_{\mathrm{loc}}({\mathbb{R}};{\mathbf{W}})$; in addition ${Y_{t}}(\eta_{0})$ is stationary.
@@
The additional characterization of ${Y_{t}}(\widehat{\mathbb{Q}}\eta_{0})$ will be useful below, when we solve the initial-value problem with a given value of $w_{0}:=\widehat{\mathbb{P}}\eta_{0}$.
@@ 2.3 Well-posedness of the stochastic process on the product space
With the characterization of the process ${Y_{t}}$ by Lemma 2.4 we can now construct the solution of the coupled microscopic system on the product space ${\mathbf{Z}}{\times}{\mathbf{H}}$ (deterministic and with finite energy) or ${\mathbf{Z}}{\times}{\mathbf{X}}$ (random and with infinite energy).
In Theorem 2.5 below we also transform the equation (2.5b) into a form that allows for a rigorous definition of a solution. The issue in (2.5b) is that the range of the operator ${\mathsf{C}}$ might not be included in the domain of the unbounded operator ${\mathbb{J}}_{\mathcal{B}}$; the integrated versions (2.8b) and (2.9b) address this.
@@
where ${Y_{t}}={Y_{t}}(\eta_{0})$. For $(z_{0},\eta_{0})\in{\mathbf{Z}}{\times}{\mathbf{H}}$, the process $t\mapsto Y_{t}(\eta_{0})$ is continuous with values in ${\mathbf{W}}$ by Lemma 2.4, and a standard ODE proof based on contraction for small times gives short-time existence and uniqueness of a solution of (2.10) with the specified regularity. The global Lipschitz bound on ${\mathrm{D}}{\mathcal{H}}_{\mathcal{A}}$ implies that the time of existence does not depend on the initial datum, and we obtain existence globally in time. Equation (2.8a) can then be recovered from (2.10) by differentiating in time, and equation (2.8b) can be used to reconstruct $\eta$. This proves the part 1 of the theorem.
@@
For our model discussed in Examples 1.1 and 1.4 we can explicitly construct the larger Hilbert space ${\mathbf{X}}\supset{\mathbf{H}}={\mathrm{L}}^{2}({\mathbb{R}};{\mathbb{R}}^{3})={\mathrm{L}}^{2}({\mathbb{R}})^{3}$. In fact, this can be done independently for each of the three components. We follow the classical construction in [Rey06] and choose
@@
Here $X$ contains distributions $T$ that can be decomposed into $T=T_{1}+T_{2}$ with $T_{1}\in{\mathrm{H}}^{-2}(R)$ and $T_{2}\in{\mathrm{L}}^{2}_{-2}({\mathbb{R}})$, i.e. $x\mapsto(1{+}x^{2})^{-1}T_{2}(x)\in{\mathrm{L}}^{2}({\mathbb{R}})$. Clearly, the shift operator $\widehat{S}_{t}:T\mapsto T(\!\;\cdot\;\!{-}t)$ generates a continuous semigroup that is the extension of the unitary shift group $(\mathrm{e}^{t\partial_{x}})_{t\in{\mathbb{R}}}$ on ${\mathrm{L}}^{2}({\mathbb{R}})$.